For the next three questions, consider the three stocks in the following table. Pt represents price at time t, and Qt represents shares outstanding at time t. Stock C splits two for one right after time 1 (P1 and Q1 represent the stock price and shares outstanding after the split) PO 0 P1 P1 1 90 100 10095 100 50 200 5 100 200 110 Question 6 (1 point) Calculate the rate of return on a price-weighted index of the three stocks for the first period (t O tot- 1). Assume the divisor value is 3. Enter your answer as a decimal, rounded to four decimal places (e.g, 0.0123) Your Answer: Answer Question 7 (1 point) If the divisor for the price-weighted index was 3 att-1, what would be the value of the divisor after split (t-1)? Round your answer to two decimal places.
Answer #1

SEE THE IMAGE. ANY DOUBTS, FEEL FREE TO ASK. THUMBS UP PLEASE

Home nert Page Layout Formulas Data Review View dd-Ins Cut E AutoSum ー E ゴWrap Text aCopy в 1 프· ー· 鱼, Δ. : rーー 逻锂函Merge & Center. $, % , 弼,8 Conditional Format eCell Insert Delete Formsat Paste Sort &Find & Format Painter Formatting as Table Styles2 Clear Clipboard Alignment Number Cells Edting A. 249 250 251 Q.6 252 253 254 255 256 257 258 259 260 261 262 Q.7 263 264 265 266 267 4 H KE CAPM UTILITY, SH PRICE WEIGHTED INDEX AT t=0, THE VALUE OF INDEX-(90 + 50 + 100)/3 = AT t-1, THE VALUE OF INDEX = (95+45+110)/3 = 80 83.33 % CHANGE OVER THE YEAR-(83.33-80)/80° 100- 4.16% INDEX AT 83.33 AT t-1 NOW SPLIT WILL NOT AFFECT THE INDEX, SO DIVISOR WILL CHANGE (95+45+ (110/2))/D 83.33 (95 45+(110/2))/83.33 D D-2.34 NSWER:B:2.34 ING PV, FV, ANNUITYDIR stru WACCRESI ex d 福 130% beta bond c t future INDEXINTL CAP BUD cleanYIELD bond 05:41 14-01-2019